

********************************************************************
UARS WEEKLY STATUS REPORT - 14 Dec 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues normal operations, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet the SOLSTICE science
objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
PEM began daylight operations on 2001333 and the PEM/AXIS-1 was also
activated on 2001333.
Verification of PACOR-A files is complete and verified. Binaries were
not the same from PACOR-A as from CDHF so these data files were processed
to science data and the science data was compared. As suspected, the version
number field contained a different value which caused the binaries to differ.
PEM performed tests of NMC data files from both the CDHF and NOAA. NMC
data files from both systems were captured and read. PEM is now capturing
NMC data directly from the NOAA web site. An automated-cron unix c-shell
script is available from PEM if it can be of use to others. For a copy,
contact: rudy@pemrac.space.swri.edu or PEM2K::RUDY.
An outstanding question from PEM is the UARS Orbit and Attitude. Some
discussion has taken place about the UARS ORBIT data but there has been no
indication on how UARS attitude data is being distributed to the
experimenters. In addition, PEM uses UARS OA Services-UOAS extensively and
that software has not been made available. It is unclear as to how attitude
information is to be determined.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 01 December 2001 (Orbit 55892, GMT Day 335) through
07 December 2001 (Orbit 55996, GMT Day 341).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 335/2003 55904 Power ON
HALOE 335/2142 55905 Power OFF
HALOE 336/1920 55919 Power ON (start)
HALOE 337/0859 55927 Power ON
PEM 337/2321 55936 AXIS OFF
WINDII 338/0101 55937 Power ON
The beta angle ranged between the angles of 16.5 degrees to 30.9 degrees
for this report period. The beta angle is now increasing towards a
maximum angle of 34 degrees on 11 December 2001 (DOY 345).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.7 to 1.0 degrees C. Additional
battery performance data for each battery on 30 November 2001 and 07
December 2001 are provided in Appendix D.
The clock error ranged from -10.6 to 13.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS 00
ALERTS 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT 24
HALOE 4
PEM 1
WINDII 1
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT 2
HALOE 1
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
* Transponder A: 00/333 - 2287.496855
* Transponder B: 01/324 - 2287.498688
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data Capture
Facility (DCF). This service should resume soon.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Another DR was written against the TAC replacement system this week.
While loading ephemeris and stored command loads a command block is
dropped in the exchange between the AP and TAC. The problem is sporadic
and has thus made it difficult to pin down. The FOT and TAC developers
will continue to investigate this issue and the other open DRs. These
issues will be resolved before the TAC is moved to full time operations.
MEETINGS
The FOT attended a UARS Eclipse Weekly on Monday, December 03, 2001.
The FOT attended a UARS Eclipse Working Meeting on Thursday, November
29, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 24 December 2001 (DOY 358).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments:
Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 07 December 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
55893 335/0243 01-704 OBC Clk Adj 0.00 to -20.6 msec/day
55897 335/0900 01-703 NBTR NOOPS D-335
55904 335/2003 01-705 HALOE Power ON
55905 335/2059 01-706 NBTR NOOPS D-336
55905 335/2142 01-707 HALOE Power OFF
55910 336/0540 01-706 NBTR PBK & NOOPS D-336
55911 336/0720 01-707 OBC Clk Adj -20.6 to 0.00 msec/day
55919 336/1920 01-708 HALOE Power ON (start)
55919 336/1920 01-709 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
55921 336/2321 01-710 NBTR NOOPS D-337
55925 337/0539 01-710 NBTR NOOPS D-337
55927 337/0859 01-711 HALOE Power ON
55935 337/2141 01-713 OBC Clk Adj 20.6 to 0.00 msec/day
55934 337/1919 01-712 NBTR NOOPS D-338
55936 337/2321 01-714 PEM AXIS OFF
55937 338/0101 01-715 WINDII Power ON
55940 338/0539 01-712 NBTR PBK & RTS-58
55941 338/0622 01-712 NBTR NOOPS D-338
55951 338-2320 01-716 NBTR NOOP 577,641
55955 339/0538 01-716 NBTR NBTR PLAYBACK
55957 339/0858 01-716 NBTR NOOPS D-339
55960 339/1300 01-716 NBTR NBTR PLAYBACK
55962 339/1601 01-717 NBTR NOOPS D-340
55970 340/0537 01-717 NBTR PBK & RTS-58
55971 340/0717 01-717 NBTR NOOP D-340
55973 340/1040 01-717 NBTR NOOPS D-340
55979 340/1917 01-718 NBTR NOOPS D-341
55986 341/0620 01-718 NBTR PBK & RTS-58
55986 341/0716 01-718 NBTR NOOPS D-341
55994 341/1916 01-719 NBTR NOOPS D-342
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 07 December 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 07 December 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
55893 335/0245 Clk Adj 0.00 to -20.6 msec/day
55911 336/0726 Clk Adj -20.6 to 0.00 msec/day
55919 336/1922 Clk Adj 0.00 to 20.6 msec/day
55935 337/2145 Clk Adj 20.6 to 0.00 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 07 December 2001
23 November 2001 - GMT Day 327 Beta = 13.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.77/-4.69 +0.0/ -0.0 10.7 0.000 0.0
2 +00.0/ -33.6 +4.48/+4.17 +28.8/-13.6 26.2 1.030 19.1
3 +22.4/ -22.4 +3.17/+2.20 +28.0/-13.6 26.2 1.030 18.6
30 November 2001 - GMT Day 334 Beta = 12.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -4.08/-4.99 +0.0/ -0.0 10.7 0.000 0.0
2 +00.0/ -33.6 +4.48/+3.17 +29.6/-13.2 26.4 1.030 19.5
3 +22.4/ -22.4 +3.17/+2.20 +28.4/-13.2 26.4 1.020 19.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 07 December 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-075 FOT 55900 340/1302 14170 Generic Late Acquisition
01-076 HALOE 55905 335/2142 N/A Haloe PWR OFF Low S/C Power
01-075 FOT 55974 340/1119 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-076 HALOE 55905 335/2142 N/A HALOE PWR OFF Low S/C Power
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-75 FOT see list Dec 01 14170 Generic Late Acquisition
SEVNTFW01/49